A woman riding a bicycle has a kinetic energy of 3600 J when traveling at a speed of 12 m/s. What is her mass?


1
Expert's answer
2021-02-05T14:06:36-0500
"KE=\dfrac{1}{2}mv^2,""m=\dfrac{2KE}{v^2}=\dfrac{2\cdot3600\ J}{(12\ \dfrac{m}{s})^2}=50\ kg."